OpenStudy (anonymous):

how do i find the limit as x approaches infinity of ( e^x/5x^2)

OpenStudy (anonymous):

you could use l'hopital's rule if it is not obvious

OpenStudy (anonymous):

you know what to do for that?

OpenStudy (anonymous):

no can you help me?

OpenStudy (anonymous):

take the derivative top and bottom separately

OpenStudy (anonymous):

don't use the quotient rule you will have to do it twice

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

\[\lim_{x\to \infty}\frac{e^x}{5x^2}=\lim_{x\to \infty}\frac{e^x}{10x}=\lim_{x\to \infty}\frac{e^x}{10}\]

OpenStudy (anonymous):

okay I'm with you so far

OpenStudy (anonymous):

you are done at that step, since that limit is \(\infty\)

OpenStudy (anonymous):

but really it should be pretty clear that \(e^x\) grows way way way faster than \(5x^2\) making the limit infinity for sure pick for example \(x=100\) the denominator is pretty big, at \(50000\) but the numerator is \(e^{100}\) which is probably too large for your calculator

OpenStudy (anonymous):

just checked with wolfram, and \(e^{100}\) has 43 zeros!

OpenStudy (anonymous):

lol thank you!